The theoretical probability of undesirable side effects resulting from taking Grebex is 2 in 19. If 570 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?
Answer by rfer(16322)

2/19=x/570
19x=570*2
19x=1140
x=60